A park is in the shape of a rectangle 8 miles long and 6 miles wide. How much shorter is your walk if you walk diagonally across the park than along the two sides?
Answer by blwinbbbles(106)

First add 6 and 8 as that is the distant your walk along the sides.
That is 14 miles...
Now you are going to solve for the distance across the rectangle...which can now actually be considered the hypotenuse of a triangle.
so a^2 + b^2 = c^2
8^2 + 6^2 = c^2
64 + 36 = c^2
100 = c^2 take the square root of both sides
10 = c
so now subtract 10 miles(the distance across)from 14 miles (the distance of the two sides) and the answer is 4 miles.
so it is 4 miles shorter to walk across the park